A block of mass M is pulled along a horizontal smooth surface by a rope of mass m. force P is applied at one end of the rope. The force which the rope exert on the block is

Grade:11

1 Answers

John Sharon Sandeep
6 years ago
Since F=ma, a=F/m. Therefore acceleration of the things=P/(M+m).Now force exerted on the object of mass M= ma =M×P/(M+m)=MP/(M+m).☺In the above step i have to find acceleration of the system in terms of only P,m,M as acc. is not given any specif